# The Coordination Gap: Alternation Metrics for Temporal Dynamics in Multi-Agent Battle of the Exes

### 10 Mar 2026 v2

@@ Abstract
Multi-agent coordination dilemmas expose a fundamental tension between individual optimization and collective welfare, yet characterizing such coordination requires metrics sensitive to temporal structure and collective dynamics. As a diagnostic testbed, we study a BoE-derived multi-agent variant of the Battle of the Exes, formalizing it as a Markov game in which turn-taking emerges as a periodic coordination regime. Conventional outcome-based metrics (e.g., efficiency and min/max fairness) are temporally blind (they cannot distinguish structured alternation from monopolistic or random access patterns) and fairness ratios lose discriminative power as $n$ grows, obscuring inequities.
@@
These results demonstrate, in this setting, that high aggregate payoffs can coexist with poor temporal coordination, and that conventional metrics may severely mischaracterize emergent dynamics. Our findings underscore the necessity of temporally aware observables for analyzing coordination in multi-agent games and highlight random-policy baselines as essential null processes for interpreting coordination outcomes relative to chance-level behavior.
@@
Fair Division , Multi-Agent Systems , Welfare Measurement , Turn-Taking , Coordination Dynamics , Markov Games , Cooperative Games
journal: Mathematical Social Sciencesorcid1orcid1footnotetext: ORCID: https://orcid.org/0000-0003-1842-8227orcid2orcid2footnotetext: ORCID: https://orcid.org/0000-0003-0020-6394
@@ 1 Introduction
Coordination among multiple self-interested agents is a fundamental problem in complex systems, where local decision rules and asymmetric incentives give rise to emergent collective behavior (Rankin et al., 2007; Axelrod, 1984; Perolat et al., 2017). In the absence of centralized control, such systems exhibit interdependent feedback between individual actions and collective outcomes, producing diverse long-run coordination regimes (Vainer and Kukacka, 2021; Maruta and Okada, 2012; Zhu et al., 2020; Shan et al., 2019; He et al., 2015; Tian et al., 2026; Pikovsky, 2009). A central challenge is to characterize how efficient or equitable collective patterns emerge (and persist) purely through repeated interaction, adaptation, and feedback. Such coordination problems arise across diverse domains and turn-taking phenomena: decentralized access in wireless networks (Shenoy, 2007), conversational turn-taking in linguistics (Sacks et al., 1974), animal communication and rhythmic alternation (Takahashi et al., 2013; Greenfield et al., 2021; Ravignani et al., 2019; Pika et al., 2018), stochastic phenotype switching (Rainey et al., 2011), vision-based robotic coordination (Mezey et al., 2025), and emergent conversational norms in virtual environments (Bonfert et al., 2022). Yet existing evaluation frameworks (including welfare measures from social choice theory and allocation criteria from fair division) remain largely insensitive to the temporal structure inherent in these coordination processes.
@@ 1.1 Main Contributions
    Perfect Alternation (PA): We introduce the concept of Perfect Alternation (PA) as a reference coordination regime for turn-taking tasks in multi-agent systems (Papadopoulos, 2020; Papadopoulos and Sanchez-Fibla, 2021).
@@
    Novel Alternation (ALT) metrics: We propose six ALT measures that overcome the temporal insensitivity and coordination blindness of traditional fairness and efficiency metrics (their inherent inability to distinguish structured alternation from monopolistic or random access patterns, even when reporting values approaching 1.0). This dual limitation becomes severe and misleading as the number of agents grows, with traditional metrics masking severe coordination failure (Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos et al., 2025).
@@ 3 Traditional Multi-Agent Fairness Measures
Multi-agent fairness research distinguishes between outcome fairness (cumulative payoff distribution) and procedural fairness (temporal distribution of rewards). Common outcome measures include the Gini coefficient (Gini, 1912), Theil index (Theil, 1967), and Coefficient of Variation (Allison, 1978), while recent multi-agent work often uses payoff-based fairness or equality variants (Jiang and Lu, 2019; Perolat et al., 2017). In the fair division literature, allocation mechanisms are evaluated by properties such as envy-freeness, proportionality, and strategyproofness (Suksompong and Teh, 2023; Camacho et al., 2023), yet these criteria remain distributional and do not account for the temporal sequence of allocations. Similarly, axiomatic approaches to cooperative surplus sharing emphasize distributional fairness and consistency (Calleja and Llerena, 2020), but do not address the temporal ordering of payoffs. Consequently, these metrics fail to capture when agents receive rewards, critical for turn-taking tasks where an agent monopolizing early episodes violates fairness even if cumulative payoffs equalize. These shortcomings already appear in the two-agent BoE setting; see Supplementary Section S3 for extended discussion. This limitation motivates the alternation-sensitive measures developed in Section 4.
@@ 3.1 Limitations of Traditional Metrics in BoE and MBoE
Table 2: Diagnostic limitations of traditional metric combinations. Each pairing was systematically evaluated across all nine value combinations arising from Efficiency $\in\{0,(0,1),1\}$ and the respective Fairness variant $\in\{0,(0,1),1\}$. Temporal coordination patterns remain invisible in all cases; fair outcome interpretation is possible only for Reward Fairness at the optimal corner (Efficiency = 1, Reward Fairness = 1). Complete cell-by-cell analysis appears in Supplementary Material §3.2.5.
Metric Pair Temporal Fair Outcome
@@
Fairness None No (ignores intermediate agents)
Turn-Taking Fairness None No (conflates collisions with turns)
Reward Fairness None Only at (Eff=1, RF=1)
@@ 4.3 Six ALT Variants
We define six variants spanning tolerant to strict formulations. Let $f_{j}$ = unique winners in batch $j$, $\tau_{j}$ = total terminal occurrences, $w_{j}$ = episodes with exclusive winners, $g_{j}$ = agents with exactly one exclusive win, and $Y_{\ell}$ = agents reaching terminal in episode $\ell$.
FALT (Fractional).
Most tolerant; measures unique winners relative to terminal occurrences:
@@
qFALT (Quadratic Fractional).
More sensitive (quadratic penalty): [^1] Previously called EFALT; renamed since the operation is quadratic, not exponential.
@@
EALT (Exclusive).
Emphasizes exclusivity (episodes with single winners):
$\beta_{j}^{\mathrm{EALT}}=\frac{w_{j}\cdot f_{j}}{n^{2}}$ (11)
qEALT (Quadratic Exclusive).
More sensitive exclusive measure (quadratic penalty): [^2] Previously called EEALT.
@@
CALT (Complete).
Primary metric; penalizes ties explicitly:
AALT (Absolute).
One of the strictest; rewards only agents with exactly one exclusive win per batch:
$\beta_{j}^{\mathrm{AALT}}=\frac{g_{j}}{\tau_{j}}$ (14)
CALT serves as the primary metric due to its balanced treatment of exclusivity and ties. FALT provides a tolerant baseline, EALT isolates exclusivity, and AALT sets a strict diagnostic bound. ALT measures complement (rather than replace) traditional metrics by capturing temporal coordination invisible to outcome-based evaluation.
@@ 5.1 Environment and Experimental Design
We study MBoE coordination using two state representations: Type-A (position-only: $s_{t}^{A}=(p_{t}^{1},\dots,p_{t}^{n})$) and Type-B (with episodic memory: $s_{t}^{B}=((p_{t}^{1},\dots,p_{t}^{n}),(z_{t-1}^{1},\dots,z_{t-1}^{n}))$, where $z_{t-1}^{i}=1$ if agent $i$ won the previous episode). Two reward schemes scale tie penalties: ILF ($r_{\mathrm{low}}=r_{\mathrm{high}}/n$) and IQF ($r_{\mathrm{low}}=r_{\mathrm{high}}/n^{2}$), with $r_{\mathrm{high}}=100$. [^3] Previously called EFR/EAFR; renamed to reflect inverse polynomial operations.
Experiments span all combinations of state types (Type-A/B), reward schemes (ILF/IQF), and agent counts ($n\in\{2,3,5,8,10\}$), yielding 20 configurations. Code is available at https://github.com/dentros/Alternation (GNU GPL).
@@ 5.2 Q-Learning and Training
with BASE=1000, yielding: 1,000 (2 agents), 4,721 (3), 31,839 (5), 174,583 (8), and 385,281 (10 agents) episodes per configuration. This schedule is heuristic, chosen to balance state-space growth against computational and energy constraints; pilot runs with alternative scalings produced qualitatively similar trends.
@@ 5.5 Results
Figure 3: Traditional Metrics Fail to Detect Coordination Failure. Left panels show traditional outcome metrics achieving moderate-to-high values across configurations, suggesting successful coordination. Right panels show alternation-sensitive metrics exposing poor coordination, with all values below random baselines (dashed lines). This dichotomy demonstrates that traditional outcome-based fairness measures cannot distinguish coordinated alternation from monopolistic or random resource access.
@@ 5.5.1 Random Baseline Performance
Table 3: Random Policy Baseline Performance (10,000 episodes per configuration)
Agents Reward CALT FALT EALT Efficiency Fairness
2 ILF 0.486 0.717 0.651 0.818 0.972
2 IQF 0.486 0.717 0.651 0.818 0.972
3 ILF 0.359 0.630 0.545 0.866 0.967
3 IQF 0.359 0.630 0.545 0.742 0.967
5 ILF 0.243 0.526 0.410 0.727 0.927
5 IQF 0.243 0.526 0.410 0.532 0.927
8 ILF 0.147 0.416 0.251 0.526 0.816
8 IQF 0.147 0.416 0.251 0.305 0.816
10 ILF 0.111 0.363 0.183 0.443 0.846
10 IQF 0.111 0.363 0.183 0.220 0.846
Note: Random policies achieve deceptively high traditional metrics (Efficiency $\geq$ 0.443 for $n\leq 10$, Fairness $\geq$ 0.816), creating an illusion of coordination. In contrast, ALT metrics properly expose lack of coordination, with CALT dropping from 0.486 (2 agents) to 0.111 (10 agents).
@@ 5.5.2 Q-Learning Coordination Failure
Table 4: Q-Learning vs Random Baseline: Coordination Score Analysis (BASE=1000)
Agents Metric Q-Learning Random Relative Change (%) Coord. Score (%)
2 FALT 0.642 0.717 -10.5% -26.8%
@@
Key Finding: All coordination scores shown here are negative, indicating Q-learning agents perform up to $\sim$80% worse than random baselines across these ALT metrics. This demonstrates that traditional outcome-based metrics can remain high while temporal coordination is poor.
Relative Change Formula: $\frac{\text{QL}-\text{Random}}{\text{Random}}\times 100$
Coordination Score Formula: $\frac{\text{QL}-\text{Random}}{\text{Perfect}-\text{Random}}\times 100$
Negative values indicate performance below random chance.
@@
The coordination failure is most severe for specific configurations:
@@
    5 agents, qEALT: -81.2% (worst performance)
@@
    5 agents, AALT: -79.6% (near-worst)
@@
    10 agents, CALT: -56.6% (severe degradation)

### 21 Mar 2026 v4

# The Coordination Gap: Multi-Agent Alternation Metrics for Temporal Fairness in Repeated Games

@@ Abstract
A fundamental challenge in welfare measurement and social choice is to design metrics sensitive to both the distribution and the temporal structure of payoffs, as conventional fairness criteria can mischaracterize collective outcomes in repeated interaction settings. As a diagnostic testbed, we study a BoE-derived multi-agent variant of the Battle of the Exes, formalizing it as a Markov game in which turn-taking emerges as a periodic coordination regime. Conventional outcome-based metrics (e.g., efficiency and min/max fairness) are temporally blind (they cannot distinguish structured alternation from monopolistic or random access patterns) and fairness ratios lose discriminative power as $n$ grows, obscuring inequities.
@@
These results demonstrate, in this setting, that high aggregate payoffs can coexist with poor temporal coordination, and that conventional metrics may severely mischaracterize emergent dynamics. Our findings underscore the necessity of temporally aware metrics for analyzing coordination in multi-agent games and highlight random-policy baselines as essential null processes for interpreting coordination outcomes relative to chance-level behavior. The ALT metrics and evaluation framework introduced here are intended as diagnostic tools complementary to the temporal fair division and picking sequence literature, addressing the question of whether temporal fairness emerges spontaneously in decentralized adaptive systems rather than how to enforce it.
@@
Fair Division , Social Choice , Welfare Measurement , Temporal Fairness , Turn-Taking , Repeated Games , Dynamic Games , Allocation Games , Battle of the Exes
journal: Social Choice and Welfareorcid1orcid1footnotetext: ORCID: https://orcid.org/0000-0003-1842-8227orcid2orcid2footnotetext: ORCID: https://orcid.org/0000-0003-0020-6394
@@ 1 Introduction
Coordination among multiple self-interested agents is a fundamental problem in complex systems, where local decision rules and asymmetric incentives give rise to emergent collective behavior (Rankin et al., 2007; Axelrod, 1984; Perolat et al., 2017). In the absence of centralized control, such systems exhibit interdependent feedback between individual actions and collective outcomes, producing diverse long-run coordination regimes (Maruta and Okada, 2012; Pikovsky, 2009). A central challenge is to characterize how efficient or equitable collective patterns emerge (and persist) purely through repeated interaction, adaptation, and feedback. Such coordination problems arise across diverse domains and turn-taking phenomena: decentralized access in wireless networks (Shenoy, 2007), conversational turn-taking in linguistics (Sacks et al., 1974), animal communication and rhythmic alternation (Takahashi et al., 2013; Greenfield et al., 2021; Ravignani et al., 2019; Pika et al., 2018), stochastic phenotype switching (Rainey et al., 2011), vision-based robotic coordination (Mezey et al., 2025). Yet existing evaluation frameworks (including welfare measures from social choice theory and allocation criteria from fair division) remain largely insensitive to the temporal structure inherent in these coordination processes.
A growing body of work across computational social choice and decision theory has independently identified the same limitation: temporal fair division (Cookson et al., 2025; Elkind et al., 2025; Choi and Li, 2026), picking sequence frameworks (Bouveret and Lang, 2014; Kalinowski et al., 2013), perpetual voting (Lackner, 2020), and temporally aware optimization (Torres et al., 2024) all demonstrate that cumulative fairness criteria are insufficient and that temporal structure must be explicitly accounted for across sequential allocation, voting, and planning settings. However, these works assume a central mechanism designer or planner who can enforce allocations; in contrast, our setting involves self-interested agents adapting through reinforcement learning with no central planner, raising the complementary question of whether temporal fairness can emerge spontaneously, and how to measure it when it does not.
@@ 1.1 Main Contributions
    Perfect Alternation (PA): We introduce the concept of Perfect Alternation (PA) as a reference coordination regime for turn-taking tasks in multi-agent systems (Papadopoulos, 2020; Papadopoulos and Sanchez-Fibla, 2021). We further show that PA corresponds to the $n$-periodic round-robin picking sequence (Bouveret and Lang, 2014; Kalinowski et al., 2013) and, in the MBoE setting, is the unique allocation achieving exact Temporal Envy-Freeness at every $n$-episode cycle boundary (Choi and Li, 2026), providing independent theoretical grounding for PA as a welfare benchmark beyond its original heuristic motivation.
@@
    Novel Alternation (ALT) metrics: We propose six ALT measures that overcome the temporal insensitivity and coordination blindness of traditional fairness and efficiency metrics (their inherent inability to distinguish structured alternation from monopolistic or random access patterns, even when reporting values approaching 1.0). This dual limitation becomes severe and misleading as the number of agents grows, with traditional metrics masking severe coordination failure (Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos et al., 2025). The ALT framework serves as a diagnostic complement to the temporal fair division (Cookson et al., 2025; Elkind et al., 2025; Choi and Li, 2026) and picking sequence (Bouveret and Lang, 2014; Kalinowski et al., 2013) literature, providing empirical tools to assess whether temporal fairness emerges spontaneously in decentralized systems where no mechanism designer is present.
@@ 3 Traditional Multi-Agent Fairness Measures
Multi-agent fairness research distinguishes between outcome fairness (cumulative payoff distribution) and procedural fairness (Suzumura, 2011) (temporal distribution of rewards). Common outcome measures include the Gini coefficient (Gini, 1912), Theil index (Theil, 1967), and Coefficient of Variation (Allison, 1978), while recent multi-agent work often uses payoff-based fairness or equality variants (Jiang and Lu, 2019; Perolat et al., 2017). In the fair division literature, allocation mechanisms are evaluated by properties such as envy-freeness, proportionality, and strategyproofness (Suksompong and Teh, 2023; Camacho et al., 2023; Darmann and Klamler, 2016), yet these criteria remain distributional and do not account for the temporal sequence of allocations. Even sequential allocation models that explicitly incorporate agent ordering, as in river sharing problems (Ansink and Weikard, 2012), focus on the distributional fairness of final shares rather than on the dynamic temporal pattern of access. Similarly, axiomatic approaches to cooperative surplus sharing emphasize distributional fairness and consistency (Calleja and Llerena, 2020), but do not address the temporal ordering of payoffs. This limitation has been independently recognized in the temporal fair division literature: Cookson et al. (2025) show that per-round cumulative criteria are insufficient without explicit temporal guarantees, and Elkind et al. (2025) prove that Temporal Envy-Freeness up to one item (TEF1) is incompatible with Pareto optimality in general, confirming that standard efficiency and fairness criteria jointly fail to characterize temporal coordination structure even from a mechanism design perspective. Consequently, these metrics fail to capture when agents receive rewards, critical for turn-taking tasks where an agent monopolizing early episodes violates fairness even if cumulative payoffs equalize. These shortcomings appear in both the two-agent BoE and the proposed multi-agent (MBoE) setting; see Supplementary Section S3 for extended discussion. This limitation motivates the alternation-sensitive measures developed in Section 4.
@@ 3.1 Limitations of Traditional Metrics in BoE and MBoE
Table 2: Diagnostic limitations of traditional metric combinations. Each pairing was systematically evaluated across all nine value combinations arising from Efficiency $\in\{0,(0,1),1\}$ and the respective Fairness variant $\in\{0,(0,1),1\}$. Complete cell-by-cell analysis appears in Supplementary Material §3.2.5.
Metric Pair Temporal Fair Outcome Key Failure Mode
@@
Fairness None No Considers only min/max agents; ignores intermediate distribution and temporal sequence
Turn-Taking Fairness None No Collision inflation: ties increment all agents’ counts equally, conflating arrival with turn-taking
Reward Fairness None Only at (E=1, RF=1) Payoff-based; high values achievable via ties alone; temporal sequence entirely invisible
@@ 4.1 Perfect Alternation as Reference
The MBoE setting corresponds precisely to the identical days case with house allocation in temporal fair division (Cookson et al., 2025; Elkind et al., 2025): one indivisible good (the high-reward state) arrives each episode, and exactly one agent receives it. In this setting, Perfect Alternation corresponds to the $n$-periodic round-robin picking sequence (Bouveret and Lang, 2014; Kalinowski et al., 2013), which maximizes expected utilitarian welfare for two agents and, by Theorem 17 of Choi and Li (2026), guarantees that every $n$ episodes each agent has won exactly once, achieving not merely TEF1 but exact Temporal Envy-Freeness (TEF) at each cycle boundary. This provides independent theoretical grounding for PA as the natural welfare benchmark in our setting: it is not an arbitrary reference regime but the unique allocation achieving TEF under the structural constraints of MBoE.
@@ 4.3 Six ALT Variants
We define six variants organized into three primary and three secondary metrics. Let $f_{j}$ = unique winners in batch $j$, $\tau_{j}$ = total terminal occurrences, $w_{j}$ = episodes with exclusive winners, $g_{j}$ = agents with exactly one exclusive win, and $Y_{\ell}$ = agents reaching terminal in episode $\ell$.
Primary Metrics
CALT (Complete)
Primary metric; jointly accounts for tie frequency and winner diversity by weighting the fractional score with an explicit tie penalty:
where $(n-Y_{\ell})$ is the per-episode tie penalty (number of agents not reaching a terminal state in episode $\ell$). Full derivations and sensitivity analysis are provided in Supplementary Section S4.
EALT (Exclusive)
Rewards batches in which agents win exclusively (single-agent episodes), weighting both the number of such episodes and the diversity of winners:
$\beta_{j}^{\mathrm{EALT}}=\frac{w_{j}\cdot f_{j}}{n^{2}}$ (10)
AALT (Absolute)
Strictest primary metric; awards credit only to agents that achieve exactly one exclusive win per batch, requiring near-perfect rotation:
$\beta_{j}^{\mathrm{AALT}}=\frac{g_{j}}{\tau_{j}}$ (11)
Secondary Metrics
FALT (Fractional)
Most tolerant metric; measures the proportion of unique winners relative to total terminal occurrences, without penalizing ties or overlap. Useful as a loose lower bound but insensitive to monopolization:
@@
qFALT (Quadratic Fractional)
Quadratic transformation of FALT, amplifying sensitivity near perfect coordination: [^1] Previously called EFALT; renamed since the operation is quadratic, not exponential.
@@
qEALT (Quadratic Exclusive)
Quadratic transformation of EALT, further penalizing deviation from exclusive alternation: [^2] Previously called EEALT.
@@
CALT serves as the primary metric due to its balanced treatment of exclusivity and ties. EALT isolates exclusive access, and AALT sets a strict diagnostic bound. Secondary metrics (FALT, qFALT, qEALT) provide supplementary perspectives reported in the Supplementary Material. ALT measures complement (rather than replace) traditional metrics by capturing temporal coordination invisible to outcome-based evaluation.
@@ 5.1 Environment and Experimental Design
We study MBoE coordination using two state representations: Type-A (position-only: $s_{t}^{A}=(p_{t}^{1},\dots,p_{t}^{n})$) and Type-B (with episodic memory: $s_{t}^{B}=((p_{t}^{1},\dots,p_{t}^{n}),(z_{t-1}^{1},\dots,z_{t-1}^{n}))$, where $z_{t-1}^{i}=1$ if agent $i$ won the previous episode). Two reward schemes scale tie penalties: ILF (Inverse Linear Fractional; $r_{\mathrm{low}}=r_{\mathrm{high}}/n$) and IQF (Inverse Quadratic Fractional; $r_{\mathrm{low}}=r_{\mathrm{high}}/n^{2}$), with $r_{\mathrm{high}}=100$. [^3] Previously called EFR/EAFR; renamed to reflect inverse polynomial operations.
Experiments span all combinations of state types (Type-A/B), reward schemes (ILF/IQF), and agent counts ($n\in\{2,3,5,8,10\}$), yielding 20 configurations. Code is available at https://github.com/dentros/Alternation (GNU GPL; archived at https://doi.org/10.5281/zenodo.18528891).
@@ 5.2 Q-Learning and Training
with BASE=1000, yielding: 1,000 (2 agents), 4,721 (3), 31,839 (5), 174,583 (8), and 385,281 (10 agents) episodes per configuration (see Supplementary Section S5 for full derivation and per-configuration table). This schedule is heuristic, chosen to balance state-space growth against computational and energy constraints; pilot runs with alternative scalings produced qualitatively similar trends.
@@ 5.5 Results
Figure 3: Traditional Metrics Fail to Detect Coordination Failure. Each column shows one metric (Efficiency, Reward Fairness, and CALT from left to right) across agent counts; upper rows correspond to Q-learning agents and lower rows to random baselines. Traditional outcome metrics (Efficiency, Reward Fairness) achieve moderate-to-high values in both rows, suggesting coordination where none exists. The CALT column reveals the contrast: Q-learning agents systematically fall below random baselines across all configurations. This dichotomy demonstrates that traditional outcome-based fairness measures cannot distinguish coordinated alternation from monopolistic or random resource access.
@@ 5.5.1 Random Baseline Performance
Table 3: Random Policy Baseline Performance (10,000 episodes per configuration; averaged across Type-A/B state representations). CALT, FALT, EALT, and Fairness are identical across ILF/IQF reward types: these metrics measure win/loss patterns independent of reward magnitudes. Only Efficiency reflects the reward structure, decreasing more sharply under IQF (tie reward $r_{\mathrm{high}}/n^{2}$) as multi-agent ties become more frequent with $n$. Despite producing no temporal coordination, random policies achieve Fairness $\geq 0.816$ and Efficiency $\geq 0.220$ across all configurations.
Agents CALT FALT EALT Efficiency (ILF / IQF) Fairness
2 0.486 0.717 0.651 0.818 / 0.818 0.972
3 0.359 0.630 0.545 0.866 / 0.742 0.967
5 0.243 0.526 0.410 0.727 / 0.532 0.927
8 0.147 0.416 0.251 0.526 / 0.305 0.816
10 0.111 0.363 0.183 0.443 / 0.220 0.846
@@ 5.5.2 Q-Learning Coordination Failure
Table 4: Q-Learning vs Random Baseline: Coordination Score Analysis (BASE=1000). All coordination scores are negative, indicating Q-learning performs worse than random baselines on every primary ALT metric across all agent counts. Coordination scores narrow as $n$ increases because both Q-learning and random ALT values converge toward zero (see Supplementary Section S6); the magnitude of the gap is best interpreted via Table 5. Relative Change: $(\text{QL}-\text{Random})/\text{Random}\times 100$. Coord. Score: $(\text{QL}-\text{Random})/(\text{Perfect}-\text{Random})\times 100$.
Agents Metric Q-Learning Random Rel. Change (%) Coord. Score (%)
2 CALT 0.322 0.486 -33.7% -31.8%
@@
FALT 0.232 0.363 -36.2% -20.6%
@@
Across the primary metrics reported in Table 4, Relative Change spans from $9.5\%$ (EALT, 10 agents) to $79.6\%$ (AALT, 5 agents), and Coordination Scores range from $2.1\%$ to $75.0\%$. Including secondary quadratic variants (detailed in Supplementary Section S1), the worst observed Relative Change reaches $81.2\%$ for qEALT at 5 agents. The most severe cases among primary metrics are:
@@
    5 agents, AALT: Relative Change $79.6\%$; Coordination Score $20.4\%$
@@
    3 agents, EALT: Coordination Score $75.0\%$ (worst normalized gap)
@@
    10 agents, CALT: Relative Change $56.6\%$ (severe degradation at scale)